The trial balance before adjustment for Grouper Company shows the following balances. Dr. Cr. Accounts Receivable $86,800 Allowance for Doubtful Accounts 2,230 Sales Revenue $462,600 Using the data above, give the journal entries required to record each of the following cases. (Each situation is independent.) 1. To obtain additional cash, Grouper factors without recourse $24,100 of accounts receivable with Stills Finance. The finance charge is 11% of the amount factored. 2. To obtain a 1-year loan of $62,900, Grouper pledges $71,900 of specific receivable accounts to Crosby Financial. The finance charge is 8% of the loan; the cash is received and the accounts turned over to Crosby Financial. 3. The company wants to maintain the Allowance for Doubtful Accounts at 7% of gross accounts receivable. 4. Based on an aging analysis, an allowance of $5,899 should be reported. Assume the allowance has a credit balance of $1,204. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when the amount is entered. Do not indent manually.)
Answer #1
1 Cash 21449
Loss on sale of receivables 2651 =24100*11%
        Accounts Receivable 24100
2 Cash 57868
Interest expense 5032 =62900*8%
       Notes payable 62900
3 Bad debts expense 8306 =(86800*7%)+2230
        Allowance for Doubtful Accounts 8306
4 Bad debts expense 4695 =5899-1204
        Allowance for Doubtful Accounts 4695